4. (20 points) Please find the even part and odd part of the following signals. a) x2(t) = sin ) b) x4(t) = e-stu(t + 2)
Answer #1

ANSWER a) Given that xey (t) = sin () =) even part of x₂ lt) is x (t) xelt) = x₂(t) + 22 (-t) - 2 Xze (t) = sin (5)+sin 1-) =Hence, odd post of 204) is mgo Lt) = sin (5) b) Xu (t) = est ult+2) =) even part of an It) is xue It) Ahelt) = 3417) + x41-